# TACA: Rethinking Cross-Modal Interaction in Multimodal Diffusion Transformers

### v2

@@ 5 Conclusion and Discussion
In this paper, we addressed two issues in MM-DiTs that limit text-image alignment in text-to-image generation: suppressed cross-attention due to token imbalance and timestep-insensitive attention weighting. We introduced Temperature-Adjusted Cross-modal Attention (TACA), a simple modification that dynamically balances multimodal interactions using temperature scaling and timestep-dependent adjustment. Combined with LoRA fine-tuning to reduce artifacts, TACA significantly improved text-image alignment on the T2I-CompBench benchmark. Our work demonstrates that strategically reweighting cross-modal interactions leads to more semantically accurate and visually coherent image generation, offering a promising approach for diffusion model research and applications.
@@ B.1 The scale of CFG guidance
Overall, the combined results across both models indicate that while increasing CFG can improve alignment to some extent, it often comes at the cost of overall performance. TACA, on the other hand, offers a more targeted and effective approach to enhancing text-image alignment, proving beneficial and complementary across different CFG scales and diffusion models.
@@ B.2 The content of the prompt
We have identified several prevalent issues regarding text alignment in state-of-the-art text-to-image models:
@@ B.3 The length of the prompt
Our proposed TACA method demonstrates comparatively more widespread effectiveness for mitigating the attribute missing issues often found in longer prompt, rather than the shorter ones. Currently, a mature benchmark for evaluating the visual-text alignment capabilities of text-to-image models with long prompts is lacking, despite the practical prevalence of longer prompts in real-world applications. Therefore, we have manually curated a set of authentic, long prompts from the internet to assess our method’s performance, and the corresponding results are presented in the following pages.
Appendix C Full parameter fine-tuning vs. LoRA
@@ Appendix D More Qualitative Results
Figure 9: Visual comparisons on visual-text alignment (FLUX.1 Dev, short prompts)
Figure 10: Visual comparisons on visual-text alignment (FLUX.1 Dev, short prompts)
Figure 11: Visual comparisons on visual-text alignment (FLUX.1 Dev, short prompts)
Figure 12: Visual comparisons on visual-text alignment (FLUX.1 Dev, short prompts)
Figure 13: Visual comparisons on visual-text alignment (SD3.5 Medium, short prompts)
Figure 14: Visual comparisons on visual-text alignment (SD3.5 Medium, short prompts)
Figure 15: Visual comparisons on visual-text alignment (FLUX.1 Dev, long prompts)
Figure 16: Visual comparisons on visual-text alignment (FLUX.1 Dev, long prompts)
Figure 17: Visual comparisons on visual-text alignment (FLUX.1 Dev, long prompts)

### v3

@@ 5 Conclusion and Discussion
In this paper, we addressed two issues in MM-DiTs that limit text-image alignment in text-to-image generation: suppressed cross-attention due to token imbalance and timestep-insensitive attention weighting. We introduced Temperature-Adjusted Cross-modal Attention (TACA), a simple modification that dynamically balances multimodal interactions using temperature scaling and timestep-dependent adjustment. Combined with LoRA fine-tuning to reduce artifacts, TACA significantly improves text-image alignment on the T2I-CompBench benchmark. Our work demonstrates that strategically reweighting cross-modal interactions leads to more semantically accurate and visually coherent image generation, offering a promising approach for diffusion model research and applications.
@@
6 Acknowledgement
This study is partially supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 (MOE-T2EP20221-0012, MOE-T2EP20223-0002), and under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ B.1 The scale of CFG guidance
Overall, the combined results across both models indicate that while increasing CFG can improve alignment to some extent, it often comes at the cost of overall performance. TACA, on the other hand, offers a more targeted and effective approach to enhancing text-image alignment, being beneficial and complementary across different CFG scales and diffusion models.
@@ B.2 The content of the prompt
We have identified several prevalent issues regarding text alignment in state-of-the-art text-to-image models. Our TACA can mitigate these issues to a certain extent.
@@ B.3 The length of the prompt
Our proposed TACA method demonstrates comparatively more widespread effectiveness for mitigating the attribute missing issues often found in longer prompt, rather than the shorter ones. Currently, a mature benchmark for evaluating the text-image alignment capabilities of text-to-image models with long prompts is lacking, despite the practical prevalence of longer prompts in real-world applications. Therefore, we have manually curated a set of authentic, long prompts from the internet to assess our method’s performance, and the corresponding results are presented in Fig. LABEL:fig:sup-7,fig:sup-8,fig:sup-9.
Appendix C Full parameter fine-tuning vs LoRA
@@ Appendix D More Qualitative Results
Figure 9: Visual comparisons on text-image alignment (FLUX.1 Dev, short prompts)
Figure 10: Visual comparisons on text-image alignment (FLUX.1 Dev, short prompts)
Figure 11: Visual comparisons on text-image alignment (FLUX.1 Dev, short prompts)
Figure 12: Visual comparisons on text-image alignment (FLUX.1 Dev, short prompts)
Figure 13: Visual comparisons on text-image alignment (SD3.5 Medium, short prompts)
Figure 14: Visual comparisons on text-image alignment (SD3.5 Medium, short prompts)
Figure 15: Visual comparisons on text-image alignment (FLUX.1 Dev, long prompts)
Figure 16: Visual comparisons on text-image alignment (FLUX.1 Dev, long prompts)
Figure 17: Visual comparisons on text-image alignment (FLUX.1 Dev, long prompts)